thatoneemochick:

i need help with math so whats 378 divide by 45 times 21 bc i keep getting the wrong answer

mhanifa:

Is it 378 : (45*21)?

thatoneemochick:

yeah i think

mhanifa:

You may use the calculator for an easy answer.

mhanifa:

Otherwise the best way is to get prime factors of 378 and 45 with 21 and simplify.

mhanifa:

378 = 2 × 3^3 × 7 45 = 3^2 x 5 21 = 3 x 7

thatoneemochick:

thank you!

mhanifa:

You are welcome!